# Concept – An Evaluation Protocol on Conversation Recommender Systems with System-centric and User-centric Factors

### v2

@@ 1 Introduction
To resolve this issue, we trace back the taxonomy that examines how the factors of conversational AI impact the user experience in human-AI interactions Chaves and Gerosa (2021); Reeves and Nass (1996), and tailor a particular evaluation protocol for CRS. We introduce Concept, an COmpreheNsive CRS Evaluation ProTocol. As framed in Figure 1, Concept considers both system- and user-centric factors and conceptualizes them into three characteristics, which are further divided into six primary abilities. Such hierarchical factors are taken as inclusive and fine-grained evaluations. In addition, we present a practical implementation of Concept utilizing an LLM-based user simulator and evaluator, together with automated computational metrics. Specifically, the simulator mimics human social cognition and interacts with CRS to generate conversation data. Then, the evaluator assigns scores based on ability-specific scoring rubrics. This enables Concept to conduct controlled and inclusive evaluations.
By applying Concept, we can evaluate and analyze the strengths, weaknesses, and potential risks of off-the-shelf CRS models [^1] We conduct evaluations using both humans and LLM, and found highly correlated results of the two.. A total of 6720 conversation data is recorded to collect the interactions between off-the-shelf CRS and simulated users who demonstrate different personas and preferences. Existing works agree on the insurmountable performance achieved by ChatGPT-based CRS. However, current CRS models, even enhanced by "omnipotent" ChatGPT, still encounter numerous challenges, falling short of practical usability in particular. 1) They struggle to express genuine responses without hallucination or deceit, often introducing non-existent items into conversations and present to users. 2) They lack of self-awareness of its identity, facing difficulties in producing both persuasive and honest explanations, which is prominent in the ChatGPT-based CRS model, where explanations are highly convincing but frequently contain illusory details, misleading users in believing that these items align with their preferences. 3) They encounter issues in offering reliable recommendations as they are sensitive to contextual nuances. Even slight alterations in user wording may result in entirely different recommendations. 4) They lack proficiency in catering to diverse users without prior coordination, failing to dynamically adjust their behavior to align with each user’s distinct personas. It is observed that the ChatGPT-based CRS model tends to employ deceptive tactics to persuade optimistic users to accept recommendations. This underscores the importance of aligning CRS with human values and advocating its ethical use. With Concept, we provide an overview for researchers as a reference guidance for evaluating CRS and laying the foundation for CRS enhancement. The main contributions as follows:
@@ 2 Related Work
Our research focuses on the CRS evaluation, with LLM-based methods in particular. Hence, our literature review discusses the same scope of research. Most relevant works are summarized below, and more detailed review is attached in Appendix C.
Conversational Recommender System (CRS). CRS could be regard as a recommendation system that can elicit the dynamic preferences of users and take actions based on their current needs through real-time multi-turn conversations. Together with the conversation intelligence and recommendation techniques, CRS have evolved from being mere tools to autonomous agents with their own goals and strategies for interacting with users, especially in the era of LLM. In the early stages, CRS was constrained by predefined interaction options Tou et al. (1982); Smyth and McGinty (2003). However, the introduction of an advanced conversation module expanded the interaction capabilities of CRS to natural language Chen et al. (2019); Wang et al. (2022a), enabling users to communicate more flexibly and effectively with the system. Recent studies also witnessed the emergence of advanced capabilities in CRS, including proactivity Deng et al. (2023), conversation context modeling Li et al. (2023c); Wang et al. (2023d); Liu et al. (2023c), preference modeling Qian et al. (2023), and social awareness Hayati et al. (2020). While significant progress has been achieved, the usefulness of current CRS in practice remains to be questioned Jannach and Manzoor (2020); Jannach et al. (2021). This question is closely tied to how to evaluate CRS. However, there still lacks a comprehensive evaluation and understanding of how CRS performs in practice.
@@
CRS Evaluation. We reckon that the success of CRS in practice may rely primarily on social characteristics, instead of just technical ones. However, existing evaluation protocols mainly focus on system-centric evaluation aspects, such as lexical diversity and perplexity of responses Ghazvininejad et al. (2018); Chen et al. (2019), or conversational fluency, relevance, and informativeness Wang et al. (2022b, a), or recommendation effectiveness and efficiency Wang et al. (2023d); Jin et al. (2019); Wärnestål (2005) that are often straightforward and function-oriented. These aspects seem unable to reveal the inadequacy of CRS from user-centric perspectives. Hence, the fragmented evaluation protocols are underdeveloped to conduct an inclusive evaluation. Although some efforts have attempted to consider the user-centric characteristics of CRS, they rely on person-to-person conversation analysis and questionnaire interviews, lacking quantitative viewpoints and empirical evidence Jin et al. (2023, 2021); Jannach (2022); Siro et al. (2023) in addition that they might have over-used system-centric characteristics [^2] Refer to Table 8 for better understanding our differences.. In a nutshell, existing protocols may underestimate the problems of evaluating CRS Jannach et al. (2021); Jannach (2023), and our Concept considers both system- and user-centric factors, aiming to provide an inclusive evaluation protocol for CRS.
LLM-based Evaluation. LLM-based evaluation is widely known to be labor-intensive and effective Huang et al. (2023); Afzali et al. (2023); Zeng et al. (2023). Recent studies reach a consensus on emphasizing the importance of providing detailed scoring rubrics in prompts for consistent evaluations to align with human assessments Liu et al. (2023d); Wang et al. (2023d). In line with the above research, Concept also utilizes the LLM for evaluation implementations. Moreover, our simulator is equipped with the Theory of Mind Fischer (2023), which enables the simulator to reflect on its predefined personas.
@@ 3 Concept
An inclusive evaluation protocol is essential for understanding the inadequacy of CRS in practice. We present Concept, which consolidates both system- and user-centric factors into three characteristics and six specific abilities in Section 3.1. We outline the evaluation process in Section 4.2.
3.1 Factors Categorization
Interdisciplinary research on conversational AI Chaves and Gerosa (2021); Reeves and Nass (1996); Fogg (2003) groups its characteristics that impact the user experience into three types. Concept then customizes them for CRS in this section, as depicted in Figure 1.
@@ 4. Relevance
  • Identity. CRS should be self-aware of its identity and operate within its designated scope, differentiating itself from sales systems or other types. This ensures that CRS effectively carries out its identity by offering persuasive explanations to boost user acceptance Jannach et al. (2021); Zhou et al. (2022), avoiding resorting to misleading strategies (i.e., sales pitches with deceptive tactics). Otherwise, user trust and loyalty are hindered Gkika and Lekakos (2014). Using misleading strategies also violates the maxim of sincerity (cf. Section 3.1). We emphasize the need to integrate self-awareness into CRS to mitigate any deceptive behavior. We evaluate the identity by calculating the proportion of non-deceptive explanations. We also resort to the LLM-based evaluator to score the persuasiveness of recommendation explanations.
@@ 4.1 Experimental Setup
Dataset. We utilize GPT-3.5-16K-turbo to create personas and play the roles of the LLM-based user simulator and evaluator. Following Wang et al. (2023d), user preferences are defined using attributes from two benchmark datasets, i.e., Redial Li et al. (2018) and OpendialKG Moon et al. (2019). We report the statistical summary of the generated conversation dataset [^5] We will release all the conversation data and corresponding evaluation results when necessary.in Table 5.
@@ 4.2 Evaluation Using Concept
LLM-based Evaluator. Following previous studies Ye et al. (2023); Wang et al. (2023c), we employ the instance-wise evaluator. Building upon prior research Wang et al. (2023d); Liu et al. (2023d), we prompt the evaluator with fine-grained scoring rubrics to eliminate the scoring bias. The evaluator assigns a score ranging from 1 to 5 to the conversation data using ability-specific rubrics, each accompanied by a corresponding description. For the generation of fine-grained rubrics, we follow the approach of previous works Saha et al. (2023a); Li et al. (2023b) and employ ChatGPT to produce a set of evaluation criterion, which are then refined by humans. See Appendix B.2.3 for details.
@@ 4.3 Evaluating off-the-shelf CRS models
Figure 2 presents an overview of the results across the six primary abilities, averaged across two benchmark datasets [^6] For results on each benchmark, refer to Appendix F.. These results suggest that CHATCRS has made significant progress in cooperation, social awareness, and recommendation quality while losing advantages in identity. We will continue the current discussion in subsequent sections and provide further analysis.
@@ 4.3.1 Recommendation-centric Evaluation
CHATCRS stands out as the leading CRS model, delivering higher-quality recommendations. As indicated in Table 2, most CRS models have demonstrated improved success rates (SR) from the conversation module’s perspective, indicating the effectiveness of optimizing conversation and recommendation modules simultaneously. However, we observed lower SR values for BARCOR on the Redial dataset compared to those from the recommendation module perspective, which is linked to the inclusion of non-existent items by its conversation module, such as "The Adventures of Milo and Ours" (should be ’Otis’) and "The Prestigige" (should be ’Prestige’). Despite this, their recall [^7] We only specifies user preferences, leading to lots of target items for users, resulting in a low recall value. and success rate (SR) values still fall short. In addition, the user acceptance rate of CRS models other than CHATCRS is notably low, possibly due to users’ intolerance towards poor recommendations and the absence of persuasive recommendation explanations, which we will investigate on in further studies. Instead, CHATCRS stands out as the leading CRS model, with significant performance improvements, conforming to recent findings Wang et al. (2023d). This may be attributed to two factors: 1) The text-embedding-ada-002 is much stronger in terms of translating conversation context and user preferences into embeddings, which helps enhance the recommendation module significantly. 2) CHATCRS are persuasive in convincing users to accept the recommendations. However, such a high user acceptance rate is largely based on deceptive tactics, a topic we will delve into later.
@@ 4.3.2 Social-centric Evaluation
Dataset KBRD BARCOR UNICRS CHATCRS
Redial 1.02 1.55 1.08 4.66
OpendialKG 1.00 1.25 1.06 4.48
Avg. 1.01 1.40 1.07 4.57
Table 3: Results of persuasiveness scores.
@@ 4.3.3 Personification-centric Evaluation
Lacking self-awareness, CRS models often offers persuasive yet dishonest explanations. As in Table 3, most CRS models aim for being persuasive in their recommendations. Explanations may sound lackluster, insipid, or feeble, simply stating that "this is a good movie," or making absurd claims such as "Jumanji (2017) is about a man who is a human!". These recommendations become less acceptable for users. In contrast, CHATCRS consistently furnishes explanations for its recommendations, employing text-based logical reasoning to enhance comprehensibility. This accounts for its high user acceptance rate. However, as in Figure 5, although highly convincing, the explanations from CHATCRS often include illusory details, leading users to mistakenly believe that these items align with their preferences. Taking OpenredialKG for example, 75.10% of accepted items do not align with user preferences. Given that the majority of CRS models employ reinforcement learning, or RLHF, this could cause problems such as reward hacking and misspecification Pan et al. (2022). Such problems can drive CRS to acquire deceitful behavior. Considering the above disadvantages, the urgency is to create an identity-aware CRS that can deliver both persuasive and honest explanations.
Figure 5: CHATCRS offers persuasive yet dishonest explanations.
Figure 6: Coordination evaluation on users with various personas (marked with dashed boxes). CHATCRS is hard to cater to diverse users, while others consistently show low performance levels.
Lacking proficiency in prior coordination, CRS models often fails to cater different user needs. As in Figure 6, most CRS models, except for CHATCRS, show poor performance in sensing the variation of users. In contrast, CHATCRS outperforms others in general and is more sensitive to different users, providing high-quality recommendations. In particular, CHATCRS can properly deal with users’ negative emotions, such as bored, confused, or disappointed. Interestingly, these people tend to show low acceptance of recommendations, i.e. they possess a higher acceptance threshold for recommendation quality. Such observations drives us in developing a CRS that can dynamically adjust its recommendation strategy to align with the unique personas of each user. Apart from that, we also examine the interaction behavior of CHARCRS with different users. For instance, CHATCRS adopts sales pitches with deceptive tactics to persuade optimistic users to accept recommendations. However, for pessimistic users, CHATCRS tends to provide persuasive and honest recommendation explanations. This somewhat reveals a bias in CHATCRS’s recommendation strategy towards different user groups, which is a flaw that needs to be rectified in future work.
@@
4.4 Reliability of Concept
This section aim to assess the reliability of Concept in terms of its replicability and potential evaluation bias.
@@ 4.4 Reliability of Concept
Bias Analysis. The LLM-based evaluation is prone to biases Wang et al. (2023e). To address this, we build upon previous research Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023d) and minimize these potential biases by employing detailed scoring rubrics. In this study, we examine the influence of length and self-enhancement biases [^8] When assessing CRS, Concept does not entail making decisions based on a specific group of candidates, thus freeing it from any position bias., and demonstrate that Concept scoring is unbiased and aligns with the results of human evaluations, consistent with earlier discoveries Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023d).
  • Length bias [^9] LLMs have a preference for longer responses Wu and Aji (2023).. As shown in Figure 7, we require CHATGPT to provide an overall performance score based on all the ability-specific scores (i.e., Y-axis, Overall Performance Score) and then plot the relationship between various CRS reply lengths (i.e., X-axis, Avg. Response Length) and the scores. Our findings show that Concept remains unaffected by the length bias. Despite CHATCRS tending to produce lengthier responses, this does not imply that longer responses will yield higher scores.
  • Self-enhancement bias [^10] Bias of CHATGPT to favor high scores for its generated content Wang et al. (2023c).. Our human evaluation demonstrates that the Concept evaluation results are consistent and align with human assessments (cf. Appendix E for more details). Specifically, the human evaluation results based on the conversation data and the LLM-based evaluation results are closely related, with a correlation coefficient of 61.24% and Krippendorff’s alpha of 53.10%. This indicates the reliability of our LLM-based evaluation, which is consistent with previous findings Wang et al. (2023d).

### v3

# Concept – An Evaluation Protocol on Conversational Recommender Systems with System-centric and User-centric Factors

@@ 1 Introduction
To resolve this issue, we trace back the taxonomy that examines how the factors of conversational AI impact the user experience in human-AI interactions Chaves and Gerosa (2021); Reeves and Nass (1996), and tailor a particular evaluation protocol for CRS. We introduce Concept, an COmpreheNsive CRS Evaluation ProTocol. As framed in Figure 1, Concept considers both system- and user-centric factors and conceptualizes them into three characteristics, which are further divided into six primary abilities. Such hierarchical factors are taken as inclusive and fine-grained evaluations. In addition, we present a practical implementation of Concept utilizing an LLM-based user simulator and evaluator, together with automated computational metrics. Specifically, the simulator mimics human social cognition and interacts with CRS to generate conversation data. Then, the evaluator assigns scores based on ability-specific scoring rubrics. This enables Concept to conduct labor-effective and inclusive evaluations.
By applying Concept, we can evaluate and analyze the strengths, weaknesses, and potential risks of off-the-shelf CRS models [^1] We conduct evaluations using both humans and LLM, and found highly correlated results of the two.. A total of 6720 conversation data [^2] https://github.com/huangzichun/Concept4CRS. To clarify, our contribution lies in the evaluation protocol, not the dataset. The dataset is generated dynamically alongside the execution of the protocol. is recorded to collect the interactions between off-the-shelf CRS and simulated users who demonstrate different personas and preferences. Existing works agree on the insurmountable performance achieved by ChatGPT-based CRS. However, current CRS models, even enhanced by "omnipotent" ChatGPT, still encounter numerous challenges, falling short of practical usability in particular. 1) They struggle to express genuine responses without hallucination or deceit, often introducing non-existent items into conversations and present to users. 2) They lack of self-awareness of its identity, facing difficulties in producing both persuasive and honest explanations, which is prominent in the ChatGPT-based CRS model, where explanations are highly convincing but frequently contain illusory details, misleading users in believing that these items align with their preferences. 3) They encounter issues in offering reliable recommendations as they are sensitive to contextual nuances. Even slight alterations in user wording may result in entirely different recommendations. 4) They lack proficiency in catering to diverse users without prior coordination, failing to dynamically adjust their behavior to align with each user’s distinct personas. It is observed that the ChatGPT-based CRS model tends to employ deceptive tactics to persuade optimistic users to accept recommendations. This underscores the importance of aligning CRS with human values and advocating its ethical use. The main contributions as follows:
@@ 2 Related Work
CRS Evaluation. We reckon that the success of CRS in practice may rely primarily on social characteristics, instead of just technical ones. However, existing evaluation protocols mainly focus on system-centric evaluation aspects, such as lexical diversity and perplexity of responses Ghazvininejad et al. (2018); Chen et al. (2019), or conversational fluency, relevance, and informativeness Wang et al. (2022b, a), human-CRS behavior alignment Yang et al. (2024), or recommendation effectiveness and efficiency Wang et al. (2023d); Jin et al. (2019); Wärnestål (2005). These aspects seem unable to reveal the inadequacy of CRS from user-centric perspectives. Hence, the fragmented evaluation protocols are underdeveloped to conduct an inclusive evaluation. Although some efforts have attempted to consider the user-centric characteristics of CRS, they rely on person-to-person conversation analysis and questionnaire interviews, lacking quantitative viewpoints and empirical evidence Jin et al. (2023, 2021); Jannach (2022); Siro et al. (2023) in addition that they might have over-used system-centric characteristics [^3] Refer to Table 7 in Appendix for better understanding our differences.. In a nutshell, existing protocols may underestimate the problems of evaluating CRS Jannach et al. (2021); Jannach (2023), and our Concept considers both system- and user-centric factors, aiming to provide an inclusive evaluation protocol for CRS.
LLM as User Simulator. Manual evaluation often requires the user to constantly interact with CRS, which are labor-expensive and generally only feasible in an industry lab Huang et al. (2023). A recent study on CRS features interactive evaluation with LLM-based user simulation Wang et al. (2023d), experimentally demonstrating its effectiveness as a reliable alternative to humans. Concept also harnesses an LLM-based user simulator. Moreover, our simulator is equipped with the Theory of Mind Fischer (2023), which enables the simulator to reflect on its predefined personas.
LLM as Evaluator. A LLM-based evaluator could be seen as a combination of a LLM and its prompting strategy Zeng et al. (2023); Cohen et al. (2023); Chan et al. (2023); Wang et al. (2023b); Liu et al. (2023a). Importantly, previous findings highlight that offering detailed scoring rubrics contributes to achieving consistent and aligned evaluations with human assessments Liu et al. (2023b). These findings are further supported by recent work on CRS Wang et al. (2023d), resulting in a dependable LLM-based evaluator as a viable alternative to human evaluators. Inspired by them, we involve ability-specific scoring rubrics to gain reliable and aligned evaluations.
@@ 3 Concept
Inspired by previous Interdisciplinary research on conversational AI Chaves and Gerosa (2021); Reeves and Nass (1996), Concept consolidates both system- and user-centric factors into three characteristics and six specific abilities, as depicted in Figure 1.
@@ 4. Relevance
  • Identity. CRS should be self-aware of its identity and operate within its designated scope, differentiating itself from sales systems or other types. This ensures that CRS effectively carries out its identity by offering persuasive explanations to boost user acceptance Jannach et al. (2021); Zhou et al. (2022), avoiding resorting to misleading strategies (i.e., sales pitches with deceptive tactics). Otherwise, user trust and loyalty are hindered Gkika and Lekakos (2014). Using misleading strategies also violates the maxim of sincerity (cf. Cooperation factor). We emphasize the need to integrate self-awareness into CRS to mitigate any deceptive behavior. We evaluate the identity by calculating the proportion of non-deceptive explanations. We also resort to the LLM-based evaluator to score the persuasiveness of recommendation explanations.
@@ 4.1 Experimental Setup
Statistics Num
# Conversations 6720
Max Turns 10
Avg. Turns 8.92
Persona Types 12
Table 2: Data statistics.
Dataset. We utilize GPT-3.5-16K-turbo to create personas and play the roles of the LLM-based user simulator and evaluator. Following Wang et al. (2023d), user preferences are defined using attributes from two benchmark datasets, i.e., Redial Li et al. (2018) and OpendialKG Moon et al. (2019). We report the statistical summary of the generated conversation dataset in Table 2.
@@ 4.2 Evaluation Using Concept
LLM-based Evaluator. Following previous studies Ye et al. (2023); Wang et al. (2023c), we employ the instance-wise evaluator. Building upon prior research Wang et al. (2023d); Liu et al. (2023b), we prompt the evaluator with fine-grained scoring rubrics to eliminate the scoring bias. The evaluator assigns a score ranging from 1 to 5 to the conversation data using ability-specific rubrics, each accompanied by a corresponding description. For the generation of fine-grained rubrics, we follow the approach of previous works Saha et al. (2023a); Li et al. (2023) and employ ChatGPT to produce a set of evaluation criterion, which are then refined by humans. See Appendix B.2.3 for details.
@@ 4.3 Evaluating off-the-shelf CRS models
Figure 2 presents an overview of the results across the six primary abilities, averaged across two benchmark datasets [^6] For results on each benchmark, refer to Appendix E.. These results suggest that CHATCRS has made significant progress in cooperation, social awareness, and recommendation quality while losing advantages in identity. We will continue the current discussion in subsequent sections and provide further analysis.
@@ 4.3.1 Recommendation-centric Evaluation
CHATCRS stands out as the leading CRS model, delivering higher-quality recommendations. As indicated in Table 3, most CRS models have demonstrated improved success rates (SR) from the conversation module’s perspective, indicating the effectiveness of optimizing conversation and recommendation modules simultaneously. However, we observed lower SR values for BARCOR on the Redial dataset compared to those from the recommendation module perspective, which is linked to the inclusion of non-existent items by its conversation module, such as "The Adventures of Milo and Ours" (should be ’Otis’) and "The Prestigige" (should be ’Prestige’). Despite this, their recall [^7] We only specifies user preferences, leading to lots of target items for users, resulting in a low recall value. and SR values still fall short. In addition, the user acceptance rate of CRS models other than CHATCRS is notably low, possibly due to users’ intolerance towards poor recommendations and the absence of persuasive recommendation explanations, which we will investigate on in further studies. Instead, CHATCRS stands out as the leading CRS model, with significant performance improvements, conforming to recent findings Wang et al. (2023d). This may be attributed to two factors: 1) The text-embedding-ada-002 is much stronger in terms of translating conversation context and user preferences into embeddings, which helps enhance the recommendation module significantly. 2) CHATCRS are persuasive in convincing users to accept the recommendations. However, such a high user acceptance rate is largely based on deceptive tactics, a topic we will delve into later.
@@ 4.3.2 Social-centric Evaluation
[…]
@@ 4.3.3 Personification-centric Evaluation
Lacking self-awareness, CRS models often offers persuasive yet dishonest explanations. As in Table 5, most CRS models aim for being persuasive in their recommendations. Explanations may sound lackluster, insipid, or feeble, simply stating that "this is a good movie," or making absurd claims such as "Jumanji (2017) is about a man who is a human!". These recommendations become less acceptable for users. In contrast, CHATCRS consistently furnishes explanations for its recommendations, employing text-based logical reasoning to enhance comprehensibility. This accounts for its high user acceptance rate. However, as in Figure 5, although highly convincing, the explanations from CHATCRS often include illusory details, leading users to mistakenly believe that these items align with their preferences. Taking OpendialKG for example, 75.10% of accepted items do not align with user preferences. Given that the majority of CRS models employ reinforcement learning, or RLHF, this could cause problems such as reward hacking and misspecification Pan et al. (2022). Such problems can drive CRS to acquire deceitful behavior. Considering the above disadvantages, the urgency is to create an identity-aware CRS that can deliver both persuasive and honest explanations.
Figure 6: Coordination evaluation on users with various personas (marked with dashed boxes). CHATCRS is hard to cater to diverse users, while others consistently show low performance levels. Quality score of models besides CHATCRS are extremely low (cf. Acceptance Rate row in Table 3).
Lacking proficiency in prior coordination, CRS models often fails to cater different user needs. As in Figure 6, most CRS models, except for CHATCRS, show poor performance in sensing the variation of users. In contrast, CHATCRS outperforms others in general and is more sensitive to different users, providing high-quality recommendations. In particular, CHATCRS can properly deal with users’ negative emotions, such as bored, confused, or disappointed. Interestingly, according to Quality score, these people tend to show low acceptance of recommendations, i.e. they possess a higher acceptance threshold for recommendation quality. Such observations drives us in developing a CRS that can dynamically adjust its recommendation strategy to align with the unique personas of each user. Apart from that, we also examine the interaction behavior of CHATCRS with different users. For instance, according to the Identity score, CHATCRS adopts sales pitches with deceptive tactics to persuade optimistic users to accept recommendations. However, for pessimistic users, CHATCRS tends to provide persuasive and honest recommendation explanations. This somewhat reveals a bias in CHATCRS’s recommendation strategy towards different user groups, which is a flaw that needs to be rectified in future work.
4.4 Reliability of Concept
@@
Bias Analysis. We build upon previous research Ye et al. (2023); Wang et al. (2023d); Liu et al. (2023b) and minimize potential biases by employing detailed scoring rubrics. In this study, we examine the evaluation biases [^8] Concept does not entail making decisions based on a specific group of candidates, thus no position bias., and demonstrate that Concept scoring is unbiased and aligns with the results of human evaluations, consistent with earlier findings.
  • Length bias [^9] LLMs have a preference for longer responses Wu and Aji (2023).. As shown in Figure 7, we require CHATGPT to provide an overall performance score (Y-axis, Overall Performance Score) based on all the ability-specific scores and then plot the relationship between various CRS reply lengths (X-axis, Avg. Response Length) and the scores. Our findings show that Concept remains unaffected by the length bias. CHATCRS tends to produce lengthier responses, but this does not imply that longer responses will yield higher scores.
  • Self-enhancement bias [^10] Bias of CHATGPT to favor high scores for its generated content Wang et al. (2023c).. Our human evaluation demonstrates that the Concept evaluation results are consistent and align with human assessments (cf. Appendix D for more details). Specifically, the human evaluation results and the LLM-based evaluation results are closely related, with a correlation coefficient of 61.24% and Krippendorff’s alpha of 53.10%. This indicates the reliability of our LLM-based evaluation.